UROP Proposal

Chris Shabsin

October 6, 1992

This fall I will be working on several projects in DOS network programming. Writing applications for DCNS will be a good way to get experience with the development process that would not be available through the classroom. There are several projects I could potentially have contact with.

One project I will be working on will be converting “lpr,” a Unix utility for remote printing, to run under DOS. This will give me a beginning look at the API for TCP/IP communications.

Other projects will include “Hesiod for Windows,” investigation of Novell’s SLIP implementation, working with Windows and Mewel API’s, and possible experimentation under alternate operating systems, like OS/2 or DesqView.